LBC Closed Dome Blue Science Tests and 3-hr Open Dome LBC checkout

Dome Closed Dome Dark HBS Off Telescope at zenith

02:50 Started series of 25 biases.

03:07 Started set of two darks, one 180 s and one 360 s. This time I remembered to close the shutter in the OB.
  • The 180s ones will be 051011 B and 051008 R. I see the wedge pattern in 051011.
  • The 360s ones will be 051341 B and 051337 R. I see this same pattern in 051341.
  • Comparing... 051341[1] avg of columns 1000:1020 is 535 for Y gt 4440 and 525 for Y lt 4440 while 051011[1] has 504 ADU for Y gt 4447 and 495 for Y lt 4447. Difference in counts is about the same for 180s and 360s darks.

03:30 Radar map looks clearer, but only one star visible.

03:34 Started set of two darks, one 10s and one 600s.
  • The 10s B dark is 053643 and the 600s B dark is 053752.
  • implot of 053643[1] columns 1000:1020 shows very slight increase at Y=4452, from 470 to 472 counts. I can barely discern this wedge pattern in the image.
  • implot of 053752[1] cols 1000:1020 shows break at Y=4444. 567 ADU below and 585 ADU above.

Another set for 60s and 900s
  • 055635 is B 60s. I can see the wedge pattern. From 473 to 478 counts at Y=4447.
  • 055824 is B 900s. Again see wedge pattern. This time the edge appears smeared, as if the pattern jumped once. From 627 to 649 ADU at 4443.
So far... for dome-dark Dark frames:
Chip 1 bias = 469 ADU in this region [1000:1020,4430:4450] yields estimates of dark count of 0.15 ADU/s for all exposure times.
exptime (s) Delta (ADU) baseline (ADU)
10 2 470
60 5 473
180 9 495
360 10 525
600 18 567
900 22 627

06:20 John went up to search for light sources
  • Found LEDs on AdSec, they will turn these off
  • LED on HBS
  • Lights in elevator room on
  • UPS on Webcam (that was turned off)
  • 3 bright LEDs on PISCES

* John and Geno going up to dome. They have some lights on. While lights are on, taking a 180s dark. 070302 B and 070256 R. Lights were turned OFF 90 sec into this exposure. On 070302, I see both wedge and bright 512-row segments/scallop patterns. implot 070302[1] 1000-1020 gives 523-531 jump at Y=4447. Delta = 8 ADU, about the same as for the dome-dark 180s exposure, although the baseline level is higher.

* Covered yellow, blue, red and green LEDs on PISCES. Turned OFF elevator room lights.

* Taking a series of dome-dark Darks. ReallyDarkSeries.ob: 180s,10s,360s,60s
  • 071337: 180s image still shows pattern. 506 to 517 at 4451. Delta=11 ADU, base 506 ADU.
  • 071707: 10s image shows a 2 ADU jump at about Y=4440. Level about 470.
  • 071748: 360s image shows a 555 to 571 jump at Y=4443. Delta=16 ADU
  • 072417: 60s image shows a jump from 481 to 487 at Y=4447.
exptime (s) Delta (ADU) baseline (ADU)
10 2 470
60 6 481
180 11 506
360 16 555

  • Dark current now higher, about 0.2 ADU/s, rate increasing with exp time.

Compare Position of Wedge vertex on Chip 3 on two 180s Darks: 051011 (910,3295) and 071337 (908,3318). Suggests pattern may shift in Y. This is what seems to have happened during the 900s dark.

07:30 Slew to Horizon so John & Geno can check LBC-Blue. They say it is completely dark.

08:00 With stairwell lights on, taking a series of 1 sec exposure at each filter (B and R), with testfiltersBR_dome.ob in LBTO Calib_OBs subdir.

08:14 Slew to zenith and repeat for 60sec exposures. Light level is low. This will also trigger tec chip so we obtain these images again.

*Next, run Matthias OB since the hard flushing was added back to the LBC code. Probably only has an effect on the time between playing the OB and the start of the first exposure.
  • Clicked Play at 08:36. Between exp and filter changes, exp about same as before.
  • Clicked Play at 08:53:01 z->V filter change 28 sec, V->z filter change 27 sec. Time between exposures 24-26 sec.

~9 UT They are opening the dome. Milky Way almost entirely visible on all-sky camera. D. Cox is called to help with opening since the new latch pins have not been used except in manual mode so far. Notes for observing:
  • 01:30, +37 is a low Galactic latitude field suitable for focal plane tilt measurements
  • "Science" OBs which test dithering and changing filters are on Triangulum galaxy (M33) which is at 1:33:50.2 30:39:36.69 and on N604 in Triangulum at 01:34:33.198 30:47:6.0. These do g-SLOAN/I-BESSEL for 60sec and SDT_Uspec/Y-FAN for 300sec. V=5.72
  • We want to test existing collimation models. New ones are in /home/okuhn/ and were copied there on 4 Sept 2011.
-rw-r--r--  1 okuhn okuhn       1899 Sep  4 01:52 SXPMLBCCollimation.dat
-rw-r--r--  1 okuhn okuhn       1761 Sep  4 01:52 DXPMLBCCollimation.dat
-rw-r--r--  1 okuhn okuhn       1854 Sep  4 05:27 SXPMLUCIFERCollimation.dat
-rw-r--r--  1 okuhn okuhn       1365 Sep  4 05:28 SXSMLUCIFERCollimation.dat

09:45 Slew to WT10_340. FPIA on it. Red was on wrong side of focus - FPIA corrections were only driving the pupil to smaller diam. Geno gives -0.5mm Global Offset. Give larger negative z. Blue went on wrong side of focus. Cleared everything and set Global Offsets = -1.5, -1.2 for SX and DX, respectively. Temp 1 (T301) and Temp 2 (glass therms) were 14.8 C and 16.9 C.

10:30 lbcrangebal on 102850, 102845. Confirmatory pair of images 103531, 103527. Good.

Focal Plane Tilt

10:35 Slew to 01:30 and +37 for star field. Elevation=78 deg. Run dofpia on field.

10:50 Start PA=0 rV superfoc. 105239/105236 ... 105510/ ~10:57 Immediately after PA=0 seq finished, ran PA=180 superfoc rV sequence.

11:00 Slew to 02:30 and 35deg to collimate. No... go to ACT0327

11:05 Loaded new collimation models. Old ones are xxxx.20110903.dat. Click initialize but hit limits. Also we have huge pointing offsets. The new collimation models were missing the 0.0 degC = Table temperature line. Added this and clicked initialize and now things look better. But in the meantime the SX mirror hit a limit. Geno raising SX now. But, on SX, Lookup table has huge offsets, -0.837 X, -233 Y, -164 Z, -137 RX, 151 RY and -66 RZ.

11:20 Copied new SX collimation model to SX.crazy.dat and SX*20110903.dat to SX*.dat

11:25 Run FPIA with new DX collimation model.

11:38 Now we have old SX and new DX collimation models. Gave -2.0mm Global Offset to SX.

11:39 Slew to WT10_310 near Triangulum galaxy. Run dofpia. Need to quit in 12 minutes... so forget range-bal and focus/collim and simply run Triangulum OB: /home/okuhn/Calib_OBs/Tests20110905/TriangulumDitherTest.ob. 114923 B, 114919 R is first pair. Does find and use guide stars. I did not check focus chip, however we did not focus beforehand. The science chip images look normal - nothing unusual in these, in particular neither of the patterns that we saw in the darks is visible at a first glance.

12:00 Closing - doing so early to give time in case of problems with doors. Few biases taken, possibly with lights on in dome.

focal plane tilt analysis of the red and blue focus sequences shows the focal planes are relatively flat. There is some difference in the focus at the edges of the Blue FP array: about 20 mic in Y and 35-40 mic in X. For the Red sequences, focus fits were not very good. But best focus measurements do not appear to differ by more then 20-30 mic from top to bottom or left to right (except for chip 1 R PA=0, and PA=180 does not confirm this).

rotator center. Compared "in-focus" images from focus sequences at PA=0,180 and find
camera Xnew Ynew PA=0 PA=180 Xold Yold
B 1037 2920 105355 105826 1035 2924
R 1078 2916 105344 105816 1078 2913

The differences are less than or equal to 4 pixels and the rotator centers written in the headers (CRPIX1, CRPIX2 for each chip) have not been changed.

14 September 2011: Results from analysis of the focal plane tilt data

Superfoc sequences at PA=0 and PA=180 were taken, and the data were analyzed using obsutil.starfocus and LBTtools.Sandbox.focfit3 Results indicate the LBC-Red is flat, and there may be a slight tilt in X for LBC-Blue, however the conditions were not so stable and the fits for the Blue PA=180 sequence, in particular, are not good. The data show almost two minima.

V PA=0

                0.013   0.005   -0.02
                4.34      4.51      4.84
 0.017                  0.014              -0.014
 4.40                     4.55                  4.81

 0.006                 -0.007              -0.011
 4.56                     4.95                 5.30
 
-0.006                  -0.011             -0.029
 4.73                     4.73                 5.31

V PA=180 degrees

                0.054   0.042   -0.02
                4.21      4.22      4.40
 0.06                   -0.0014             -0.036
 3.95                     4.32                  4.14

 0.006                 -0.042              -0.044
 4.34                     4.08                  4.24
 
 0.032                  -0.038             -0.040
 4.53                     4.21                  4.27

r PA=0 degrees

                -0.011   -0.0052   -0.01
                 5.87       5.65        5.86
 -0.019               -0.024             -0.0027
  5.39                   5.33                6.12

 -0.024                 -0.008               0.04
  5.48                     5.39                 6.02
 
 0.0                      0.013               0.061
 5.47                      5.25                5.87

r PA=180 degrees

            -----       -----          -----
            -----       -----          -----       
 0.009                -----             0.0122
  4.90                 -----               5.55

 0.013                -----              0.0156
  5.13                 -----                5.76
 
 0.016                -----              0.024
  5.80                 -----                6.16

-- OlgaKuhn - 07 Sep 2011

  • V PA=0:
    focus_V0.jpg

  • V PA=180:
    focus_V180_3x4.jpg

  • r PA=0:
    focus_r0.jpg

  • r PA=180:
    focus_r180.jpg
I Attachment Action Size Date Who Comment
focus_V0.jpgjpg focus_V0.jpg manage 52 K 14 Sep 2011 - 19:00 OlgaKuhn V PA=0
focus_V180.jpgjpg focus_V180.jpg manage 52 K 14 Sep 2011 - 19:00 OlgaKuhn  
focus_V180_3x4.jpgjpg focus_V180_3x4.jpg manage 53 K 14 Sep 2011 - 19:01 OlgaKuhn V PA=180
focus_V180all.jpgjpg focus_V180all.jpg manage 55 K 14 Sep 2011 - 19:01 OlgaKuhn  
focus_r0.jpgjpg focus_r0.jpg manage 52 K 14 Sep 2011 - 19:00 OlgaKuhn r PA=0
focus_r180.jpgjpg focus_r180.jpg manage 52 K 14 Sep 2011 - 19:01 OlgaKuhn r PA=180
Topic revision: r5 - 14 Sep 2011, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback